# Impact of 5G SA Logical Vulnerabilities on UAV Communications: Threat Models and Testbed Evaluation

### 04 Mar 2026 v1

@@ IV-B Threat Model 2: Insider Threat
Figure 2: Threat Model 2: Insider Adversary with Access to the 5G Core Network.
@@
After the association is established with the UPF, the script initiates the following phases: (i) iteration over a range of SEIDs passed as parameters (brute-force enumeration); (ii) for each SEID, a Session_Deletion_Request with a unique sqn is sent and the deletion response is analyzed; and (iii) when the response indicates Cause = Request Accepted, the corresponding PDU Session is terminated.
The result observed in the experiments is the teardown of the PDU Session (and its GTP-U tunnel) for each SEID for which the UPF returned Request Accepted. Regardless of whether only the UAV or GCS session was interrupted, the observed behavior was the UAV entering failsafe mode and returning to its point of origin, impacting its operations.
@@
IV-C Threat Model 3: Compromised gNodeB
@@
This threat model considers an adversary that has gained control of a gNodeB (gNB) that is already integrated into the legitimate 5G Standalone network. Unlike an external fake base station, the compromised gNB remains part of the architecture: it participates in the Service-Based Architecture (SBA) and in the user plane (GTP-U), so it can observe and manipulate both control-plane signalling (N2/NAS) and the Command and Control (C2) traffic of UEs attached to it. Compromise may occur through realistic vectors such as exploitation of known vulnerabilities (CVEs) in virtualized RAN or O-RAN components, unauthorized access to management interfaces (e.g., SSH, NETCONF, O1/OAM), supply-chain attacks in cloud-native deployments, or failures of isolation between network functions in shared infrastructure [4]. In private 5G networks used for critical operations (e.g., BVLOS UAV missions), the larger attack surface of the access node makes a logically compromised gNB a plausible and relevant internal threat [7].
Why the gNB is a critical point. The gNB is the termination point of the radio link ciphering (PDCP). Even when the air interface is protected (e.g., NEA2), the gNB must decapsulate and decrypt user-plane packets to forward them toward the core. Thus, inside a legitimate gNB, user data (including the inner IP and application payload) is available in cleartext. The 3GPP specifications do not mandate encryption or integrity protection on the N3 interface (GTP-U between gNB and UPF): protection is often left to operator choice, and in practice many deployments do not use IPsec or integrity on N3 [9]. Operators are reluctant to deploy IPsec on the N3 interface because it is CPU-intensive and reduces the throughput of user traffic; moreover, user data is often perceived to be protected at the application layer (e.g., TLS), leading some to consider IPsec redundant [2]. While 3GPP recommends securing the N2 (NGAP) and N3 (GTP-U) interfaces using IPSec, encryption is not mandatory, only integrity protection is strongly advised, leaving these interfaces accessible for monitoring in many deployments [5]. Studies report that GTP-capable nodes, including UPFs, are in some cases exposed to the public Internet, contradicting the ”walled garden” assumption: measurements have identified hundreds of thousands of valid GTP hosts accessible via the public Internet across multiple countries and providers [11]. For the user plane on the radio and backhaul segments, integrity protection is frequently optional (e.g., for fronthaul and N3): integrity protection of fronthaul traffic is optional in the protocol standards due to performance requirements of the eCPRI protocol, and lack of integrity over fronthaul is deemed acceptable by some specifications under perceived low risk [10]; empirical studies of live deployments show that User Plane integrity (e.g., NIA2) is often absent, with only a minority of operators enabling it [3, 4]. Together, this increases the risk that a compromised gNB can read and modify traffic without the network or the UE detecting it. In summary: (1) the gNB terminates radio encryption and thus has cleartext access to the payload; (2) the N3/GTP-U path often has no cryptographic integrity; (3) a compromised gNB can therefore tamper with user-plane traffic, including C2 protocols carried inside the tunnel.
Attack: UAV Navigation Hijacking via Compromised gNB Data Tampering (ATK-ROGUE_GNB-01). We consider the case where the UAV is served by a legitimate gNB (gNB01) and the GCS is served by a compromised gNB (gNB-rogue). All C2 traffic from the GCS to the UAV traverses the 5G core and, on the return path to the UAV, may pass through the compromised gNB if the topology so dictates; in our testbed we arrange so that the GCS is attached to the rogue gNB, so that uplink C2 (GCS → UAV) crosses the compromised node. The attacker’s goal is to alter navigation commands without breaking the tunnel or the radio link. MAVLink messages such as SET_POSITION_TARGET_LOCAL_NED carry waypoint or position targets; they are transported in cleartext inside the GTP-U tunnel. A compromised gNB that can inspect and modify GTP-U payloads can replace the target coordinates with values chosen by the attacker, causing the UAV to fly toward an unintended location while the operator believes the original command was executed. This corresponds to MITRE ATT&CK Impact (TA0040) and the technique T1565 (Data Manipulation).
@@
To emulate this model we use a custom script, which deploys a 5G SA environment in a Kubernetes (Kind) cluster with Open5GS (AMF, SMF, UPF, etc.) and security hardening on the AMF (NIA0/NEA0 removed). The user plane is left without N3 encryption/integrity, consistent with 3GPP optionality and common practice. There are three UERANSIM gNBs: gNB01 (nci 0x10), gNB02 (nci 0x20, no UEs), and gNB-rogue (nci 0x30), the latter under attacker control but still connected to the legitimate core. The UAV (IMSI 001) attaches only to gNB01 and runs a MAVLink simulator; the GCS (IMSI 002) and a ROGUE UE (IMSI 003) attach to gNB-rogue. Thus, all GCS traffic, including C2 toward the UAV, passes through the compromised gNB. PLMN is 999/70 and a single slice (SST 1, SD 0x111111) is used.
[…]
@@
Observed result. In the testbed, when the GCS sends a position target, the modified command causes the UAV to move toward the attacker’s chosen coordinates. The operator sees the command accepted by the UAV but not the altered target, demonstrating navigation hijacking via data manipulation. The vulnerability is cross-layer: the 5G user plane often does not provide integrity for N3/GTP-U, and the C2 application (MAVLink) relies on that channel without end-to-end integrity (e.g., signing). A compromised gNB exploits both aspects.
@@
Mitigation. Mitigations align with the evidence from the literature and slides: (i) User-plane integrity: where feasible, enable integrity (and optionally confidentiality) on the N3 segment (e.g., IPsec between gNB and UPF) or future 3GPP mechanisms for user-plane integrity, so that tampering by a compromised node becomes detectable. (ii) Application-layer protection: use MAVLink signing (or equivalent integrity/authenticity) so that the UAV verifies origin and integrity of C2 commands before execution; this protects even if the 5G path is compromised. (iii) Monitoring and segmentation: monitor C2 traffic for anomalies at gNB/UPF and isolate C2 in a dedicated slice or session with stricter policies, reducing the impact of a single compromised gNB.

### 12 Mar 2026 v2

@@ III-B Threat Model 2: Insider Threat
Figure 4: Threat Model 2 - Insider Adversary with Access to the 5G Core Network.
@@
After the PFCP association is established with the UPF, the script initiates the following phases: (i) iterating over a range of SEIDs provided as parameters (brute-force enumeration); (ii) for each SEID, a PFCP Session Deletion Request with a unique sequence number is sent and the corresponding response is analyzed; and (iii) when the response indicates the cause value “Request accepted”, the corresponding PDU Session is terminated. An important implementation detail is that UPF terminates the association with a peer that does not respond to PFCP Heartbeat Requests. Therefore, one of the distinguishing features of the algorithm that allowed SMF spoofing, unlike other studies that execute the attack from within SMF itself, for example [7], is that the script responds with a Heartbeat Response at the right time so that the UPF does not disassociate the “insider” and the PFCP session remains active. The Listing 4 shows the logs from UPF when the Session Deletion Request is successful.
Listing 4: logs of the UPF with session removed.
[upf] INFO: [Removed] Number of UPF-sessions is now 2 (../src/upf/context.c:253)
The result observed in the experiments is the teardown of the PDU Session (and its GTP-U tunnel) for each SEID for which the UPF returned Request Accepted. Regardless of whether only the UAV or GCS session was interrupted, the observed behavior was the UAV entering failsafe mode and returning to its point of origin, impacting its operations. The Listing 5 shows the UAV’s real-time log, where after the GCS lost its session (and its respective GTP-U tunnel) and the UAV could no longer communicate with it, it entered failsafe mode and returned to HOME to perform the landing, causing an interruption in its operations.
Listing 5: logs of the UAV state after the attack (forced landing).
... (repeated)
[UAV] heartbeat sent
[UAV] FAILSAFE: GCS heartbeat lost (>=60s). Entering RTL.
[UAV] RTL: returning to HOME.
[UAV] trying to restore link (waiting for GCS heartbeat)...
[UAV] heartbeat sent
... (repeated)
[UAV] heartbeat sent
[UAV] RTL: reached HOME, starting landing.
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] landed.
[UAV] trying to restore link (waiting for GCS heartbeat)...
[UAV] heartbeat sent
... (repeated)
III-C Threat Model 3: Compromised gNodeB
@@
The main papers on attacks involving the gNodeB focus on fake base station (fake gNodeB) attacks [42], [32], [23]. In our final threat model, we consider a different approach: a compromised gNodeB. Unlike an external fake base station, the compromised gNodeB remains part of the architecture: it participates in the SBA and in the User Plane (GTP-U), so it can observe and manipulate both control-plane signaling (N2/NAS) and the C2 traffic of UEs attached to it. Compromise may occur through realistic vectors such as exploitation of known vulnerabilities (CVEs) in virtualized RAN or O-RAN components, unauthorized access to management interfaces (e.g., SSH, NETCONF, O1/OAM), supply-chain attacks in cloud-native deployments, or failures in isolation between network functions in shared infrastructure [21]. In private 5G networks used for critical operations (e.g., BVLOS UAV missions), the larger attack surface of the access node makes a logically compromised gNodeB a plausible and relevant internal threat [32]. Furthermore, recent military events such as the Russo-Ukrainian War illustrate how mobile communication has become part of the battlefield, where the physical compromise of telecommunications assets makes this scenario more realistic [13]. In short, the difference between compromising a gNodeB and creating a fake gNodeB is that it is a trusted element of the 5G network, potentially creating additional vulnerability and exploitation vectors.
The gNodeB is a critical point in the architecture because it terminates the radio link encryption (PDCP). Even with protection on the air interface (e.g., NEA2), gNodeB needs to decapsulate and decrypt the User Plane packets to forward them to the 5GC. Therefore, within the legitimate gNodeB, user data may be available in plain text depending on the encryption used at the application layer. The 3GPP specifications do not mandate encryption or integrity protection on the N3 interface (GTP-U between gNodeB and UPF) [5]. As previously mentioned, protection is generally at the operator’s discretion, and in practice, many deployments do not use IPsec or integrity protection on N3 [21][24][38]. One of the reasons operators are reluctant to deploy IPsec on the N3 interface is because it is CPU-intensive and reduces the throughput of user traffic. Moreover, user data is often perceived to be protected at the application layer (e.g., TLS), leading some to consider IPsec redundant.
The 3GPP specifications recommend securing N2 (NGAP) and N3 (GTP-U) interfaces with IPSec, for example, but encryption is not mandatory, and and only integrity protection is strongly recommended. As an example of deployment flaws, numerous GTP-compliant nodes, including UPFs, have been found directly exposed to the Internet, in large numbers and across various countries and operators, contradicting the “walled garden” assumption of 5G. In the User Plane, integrity is considered optional because. For example, in the fronthaul, integrity protection is frequently avoided due to eCPRI performance constraints in the N3 interface. studies indicate that integrity in the User Plane is used by a small number of operators [21][41].
@@
To evaluate this scenario in our testbed, we developed a custom deployment script that instantiates a 5G SA environment in a Kubernetes (Kind) cluster using Open5GS (AMF, SMF, UPF, etc.). User Plane traffic over N3 is left without encryption or integrity protection, consistent with the optional nature of User Plane protection in 3GPP specifications. The deployment includes three UERANSIM gNodeB instances: gNodeB01 (nci 0x10) and gNodeB-rogue (nci 0x30), the latter being under attacker control while remaining connected to the legitimate core network. The UAV (IMSI 001) attaches only to gNodeB01 and runs a MAVLink-based UAV simulator; the GCS (IMSI 002) and a ROGUE UE (IMSI 003) associate with gNodeB-rogue (Figure5). Consequently, all GCS traffic, including C2 toward the UAV, passes through the compromised gNodeB. PLMN is 999/70 and a single slice (SST 1, SD 0x111111) is used.
Figure 5: Threat Model 3 - Compromised gNodeB.
In this scenario, the attacker’s objective is to alter the UAV’s operation without disrupting the logical tunnel or radio link connection. MAVLink messages such as SET_POSITION_TARGET_LOCAL_NED carry waypoint or position commands; they are transmitted in cleartext inside the GTP-U tunnel, even when MAVLink 2.0 is used, as in our testbed. With access to the compromised gNodeB, the attacker can inspect and modify the GTP-U payloads, which can reposition, for example, the UAV to another location desired by the threat, while the legitimate operator (GCS) believes that the UAV is executing the original command.
Figure 6: Attack sequence on compromised gNodeB.
In our experiment, we assume that the attacker has gained root access to the compromised gNodeB shell. With this access, the attacker can install tools such as iptables, Python, and Scapy, etc. Using iptables, traffic on port 2152 (GTP-U) is redirected to a userspace queue NFQUEUE. The script used to manipulate the packets works as follows: (i) decapsulates the GTP-U payload to obtain the inner IP packet (UE-UPF traffic); (ii) parses the inner UDP payload (e.g., MAVLink on ports 14550/14551); (iii) updates the current UAV position from GLOBAL_POSITION_INT when present; (iv) on SET_POSITION_TARGET_LOCAL_NED, replaces the target with an offset from the current position toward attacker-defined coordinates (e.g., MAV_FRAME_LOCAL_OFFSET_NED); (v) re-encapsulates and re-injects the packet. No radio or application-layer encryption is broken; the gNodeB already has cleartext access to the tunnel payload. Algorithm 3 summarizes the data path and tampering step.
The attacker maintains an estimate of the UAV’s current position from MAVLink telemetry traversing the gNB. Whenever a GLOBAL_POSITION_INT message is observed in the GTP-U stream (e.g., on the downlink toward the GCS), the script extracts latitude, longitude, and relative altitude:
$\begin{split}\phi_{\mathrm{uav}}&=\frac{\mathtt{lat}}{10^{7}},\quad\lambda_{\mathrm{uav}}=\frac{\mathtt{lon}}{10^{7}},\\ h_{\mathrm{uav}}&=\frac{\mathtt{relative\_alt}}{1000}\quad\text{(m)}.\end{split}$ (1)
The attacker’s desired destination is given in geographic coordinates and altitude $(\phi_{\mathrm{rogue}},\lambda_{\mathrm{rogue}},h_{\mathrm{rogue}})$. To inject a command that the autopilot interprets as “move from current position,” the script uses the frame MAV_FRAME_LOCAL_OFFSET_NED (value 7), in which the fields $x$, $y$, and $z$ are offsets in meters in the North–East–Down (NED) frame centered on the UAV’s current position.
The conversion from geographic difference to NED meters uses a planar approximation (valid at typical BVLOS scales). Latitude yields approximately 111 320 m per degree; longitude depends on latitude:
$\begin{split}\Delta N&=(\phi_{\mathrm{rogue}}-\phi_{\mathrm{uav}})\times 111\,320,\\ \Delta E&=(\lambda_{\mathrm{rogue}}-\lambda_{\mathrm{uav}})\times 111\,320\\ &\quad\times\max(0.2,\,\cos\phi_{\mathrm{uav}}),\end{split}$ (2)
with angles in degrees and $\phi_{\mathrm{uav}}$ in radians in the cosine term. The vertical offset in NED ($z$ positive down) is
$\Delta z=-(h_{\mathrm{rogue}}-h_{\mathrm{uav}}).$ (3)
The script replaces in the intercepted SET_POSITION_TARGET_LOCAL_NED the fields coordinate_frame =7=7, x $\Delta N$, y $\Delta E$, z $\Delta z$, so that the UAV interprets “move $\Delta N$ m north, $\Delta E$ m east, and $\Delta z$ m in altitude from current position,” effectively reaching $(\phi_{\mathrm{rogue}},\lambda_{\mathrm{rogue}},h_{\mathrm{rogue}})$. The attacker therefore does not need to know the drone’s HOME (black-box context); observing telemetry to update $(\phi_{\mathrm{uav}},\lambda_{\mathrm{uav}},h_{\mathrm{uav}})$ and recomputing the offset for each intercepted command is sufficient.
@@
During testing on our testbed, we observed that when the GCS sends a position update command, the UAV moves toward coordinates defined by the attacker, after manipulation at the compromised gNodeB. The operator observe that the command was accepted by the UAV but not observe the altered navigation target, demonstrating navigation hijacking via data manipulation. The vulnerability is cross-layer: the 5G User Plane often does not provide integrity for N3/GTP-U, and the C2 application (MAVLink) relies on that channel without end-to-end integrity (e.g., signing). The compromised gNodeB exploits both aspects. The Listings LABEL:lst:gnb-rogue-exploit and 6 shows the logs from the attackers perspective and the maniputaled command accepted by the UAV.
Algorithm 3 UAV Navigation Hijacking (Rogue gNB; offset from current position)
1:procedure RogueGnB-SetposHijack-Offset
2:  $ROGUE\_LAT,ROGUE\_LON,ROGUE\_ALT\leftarrow\text{attacker destination}$; $QUEUE\_NUM\leftarrow 1$; $GTP\_U\_PORT\leftarrow 2152$
3:  $_{u}av\_lat,_{u}av\_lon,_{u}av\_alt\_m\leftarrow\text{None} \triangleright$ UAV position from telemetry
4:  BindNFQUEUE(QUEUE_NUMQUEUE\_NUM, modify_packet)
5:  loop (each packet pktpkt)
6:   if $\text{port}(pkt)\neq GTP\_U\_PORT$ then Accept(pktpkt); continue
7:   $header,optional,innerBytes\leftarrow\textsc{ExtractGtpUInner}(pkt)$
8:   if $innerBytes=\emptyset$ then Accept(pktpkt); continue
9:   $msgs\leftarrow\textsc{ParseMAVLink}(innerBytes)$
10:   for each $msg\in msgs$ do
11:    if $msg.\text{type}=\text{GLOBAL\_POSITION\_INT}$ then
12:     $_{u}av\_lat,_{u}av\_lon,_{u}av\_alt\_m\leftarrow msg.\text{lat}/10^{7},msg.\text{lon}/10^{7},msg.\text{relative\_alt}/1000$
13:   for each $msg\in msgs$ do
14:    if $msg.\text{type}=\text{SET\_POSITION\_TARGET\_LOCAL\_NED}\wedge_{u}av\_lat\neq\text{None}$ then
15:     $north\_m,east\_m\leftarrow\textsc{LatLonDeltaToNED}(ROGUE\_LAT,ROGUE\_LON,_{u}av\_lat,_{u}av\_lon)$
16:     $z\_ned\leftarrow-(\text{ROGUE\_ALT}-_{u}av\_alt\_m)$
17:     $msg.\text{coordinate\_frame}\leftarrow\text{MAV\_FRAME\_LOCAL\_OFFSET\_NED}$; $msg.x\leftarrow north\_m$; $msg.y\leftarrow east\_m$; $msg.z\leftarrow z\_ned$
18:     $newInner\leftarrow\textsc{RebuildIP}(msgs)$; $newGtp\leftarrow\textsc{GtpUBuild}(header,optional,newInner)$
19:     SetPayload(pktpkt, newGtpnewGtp); break
20:   Accept(pktpkt)
21:end procedure
@@
To mitigate this issue, we propose solutions aligned with findings reported in the literature: (i) User Plane integrity: where feasible, enable confidentiality and integrity protection on the N3 segment (e.g., IPsec between gNB and UPF) or future 3GPP mechanisms for User Plane integrity, so that tampering by a compromised node becomes detectable. (ii) Application-layer protection: use MAVLink signing (or equivalent integrity/authenticity) so that the UAV verifies origin and integrity of C2 commands before execution; this protects even if the 5G path is compromised. (iii) Monitoring and segmentation: monitor C2 traffic for anomalies at gNB/UPF and isolate C2 in a dedicated slice or session with stricter policies, reducing the impact of a single compromised gNB.
Listing 6: Logs of the UAV.
... (repeated)
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] received: SET_POSITION_TARGET_LOCAL_NED | content: {’mavpackettype’: ’SET_POSITION_TARGET_LOCAL_NED’, ’time_boot_ms’: 3294353477, ’target_system’: 1, ’target_component’: 0, ’coordinate_frame’: 7, ’type_mask’: 4039, ’x’: -1424.2503662109375, ’y’: 2380.08154296875, ’z’: -50.0, ’vx’: 0.0, ’vy’: 0.0, ’vz’: 0.0, ’afx’: 0.0, ’afy’: 0.0, ’afz’: 0.0, ’yaw’: 0.0, ’yaw_rate’: 0.0}
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
... (repeated)
III-D Architectural Weaknesses and Security Implications
The attacks presented in this work rely on architectural weaknesses and deployment assumptions in the interaction between the 5G architecture and UAV communication protocols. They do not depend on implementation bugs or software vulnerabilities. The absence or optional use of security mechanisms on certain interfaces (e.g., N3 and N4) and the lack of application-layer authentication in MAVLink when message signing is disabled allow command injection, session disruption, and navigation manipulation.
To characterize these issues from a security engineering perspective, we map the evaluated attacks to relevant classes from the Common Weakness Enumeration (CWE) [37]. The mapping shows that the demonstrated attacks fall under known categories of architectural weaknesses, such as insufficient authentication, improper access control, or lack of integrity protection.
TABLE I: Summary of threat models, attacks, UAV impact, mitigations, and related CWEs.
Threat Model Attack Exploited Interface UAV Impact Proposed Mitigation Related CWE
Rogue UE MAVLink command injection UE in same slice/DNN Forced landing / unauthorized commands MAVLink signing CWE-345: Insufficient Verification of Data Authenticity; CWE-306: Missing Authentication for Critical Function
Insider PFCP session deletion N4 (SMF–UPF) UAV failsafe due to session loss N4 integrity protection CWE-284: Improper Access Control; CWE-862: Missing Authorization
Compromised gNodeB MAVLink payload manipulation N3 (GTP-U) Navigation hijacking N3 integrity + MAVLink signing CWE-345: Insufficient Verification of Data Authenticity; CWE-319: Cleartext Transmission of Sensitive Information
Table  I summarizes the evaluated threat models, the exploited architectural weaknesses, the affected 5G interfaces, and the resulting impact on UAV command and control. It also lists the corresponding CWE categories and mitigation strategies.
This analysis indicates that the demonstrated attacks are manifestations of architectural weaknesses that may affect UAV systems relying on cellular connectivity.
IV Related Work
With the aim of analyzing related work, we group prior work by where the vulnerability lies (UE, RAN, core, User Plane, Control Plane) and by general drone security research, then compare each category to our threat models and testbed.
UE vulnerabilities. Lisowski et al. (2024) [22] treat SIM cards as an attack vector and show that compromised identity modules can affect devices and basebands. We do not rely on identity compromise. In Threat Model 1 (TM1) a rogue UE is already on the same slice and DNN as the UAV; the vulnerability lies at the application layer. We show that without application-level authentication, an attacker can inject MAVLink C2 and force the UAV to land. Olimid and Nencioni (2020) [27] survey 5G slice security and point to weak isolation as an enabler of attacks. Our TM1 turns that into a concrete scenario: one rogue UE in the same logical network can reach the UAV and send commands that the autopilot accepts.
RAN vulnerabilities. Bennett et al. (2024) [11] fuzz RAN-core interfaces such as N2 to find protocol bugs in base stations and cores. We do not look for new RAN bugs. In Threat Model 3 (TM3), the gNodeB is already compromised and still trusted by the core. We show that it can alter User Plane traffic on the N3 path. Because MAVLink runs in cleartext inside GTP-U, we tamper with C2 messages (e.g., SET_POSITION_TARGET_LOCAL_NED) and demonstrate navigation hijacking while the link appears normal to the operator. Wen et al. (2024) [40] offer detection of L3 attacks from O-RAN telemetry. We focus on how a compromised RAN node can mount such an attack, not on detection.
Core vulnerabilities. Sun et al. (2025) [35] and Dong et al. (2025) [17] fuzz the 5G core to find stateful or logical bugs. In Threat Model 2 (TM2), we assume an insider who can reach N4; we do not discover new core flaws; we use known PFCP behavior. Our attacker sets up a PFCP association with the UPF, sends Session Deletion Requests, keeps the association alive with heartbeats, and we measure how UAV-GCS connectivity and failsafe behave when User Plane are torn down. Amponis et al. (2022) [7] also use PFCP from inside the core and evaluate the effect on drone swarms. We go further by having the insider spoof an SMF peer and by giving testbed results for a single UAV-GCS link and failsafe.
User Plane vulnerabilities. Zhang et al. (2025) [43] measure GTP in the wild and find many nodes and message types exposed and usable for denial-of-service and session hijacking. In TM3, we rely on the fact that N3 often lacks integrity protection: a compromised gNodeB sees and can change GTP-U payloads, so it can rewrite MAVLink commands before they reach the UAV. Xing et al. (2024) [41] show that the lack of integrity on the fronthaul allows software-based attacks at scale. We do not touch the fronthaul; we show that the same idea applies to N3 and that without integrity or application-layer signing, C2 can be tampered with by a compromised gNodeB.
TABLE II: Comparison between this work and related literature. $\bullet$: functional presence; $\circ$: absence; $\diamond$: limited or partial presence. This work (last row) addresses all seven criteria.
Reference UE RAN Core N3/GTP-U N4/Control Plane UAV Testbed
Lisowski et al. (2024) [22] $\bullet$ $\circ$ $\circ$ $\circ$ $\diamond$ $\circ$ $\bullet$
Olimid, Nencioni (2020) [27] $\diamond$ $\circ$ $\diamond$ $\circ$ $\circ$ $\circ$ $\circ$
Bennett et al. (2024) [11] $\diamond$ $\bullet$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$
Wen et al. (2024) [40] $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$ $\bullet$
Sun et al. (2025) [35] $\diamond$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$
Dong et al. (2025) [17] $\diamond$ $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$
Amponis et al. (2022) [7] $\circ$ $\circ$ $\bullet$ $\diamond$ $\bullet$ $\bullet$ $\bullet$
Zhang et al. (2025) [43] $\circ$ $\circ$ $\diamond$ $\bullet$ $\circ$ $\circ$ $\bullet$
Xing et al. (2024) [41] $\circ$ $\bullet$ $\circ$ $\diamond$ $\circ$ $\circ$ $\bullet$
Cremers, Dehnel-Wild (2019) [15] $\diamond$ $\circ$ $\diamond$ $\circ$ $\bullet$ $\circ$ $\circ$
Salazar et al. (2021) [31] $\circ$ $\diamond$ $\bullet$ $\diamond$ $\bullet$ $\circ$ $\bullet$
Nassi et al. (2021) [26] $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
Branco et al. (2025) [12] $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
Du et al. (2024) [18] $\diamond$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$
Veksler et al. (2024) [39] $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\bullet$ $\bullet$
This work $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
Control Plane vulnerabilities. Cremers and Dehnel-Wild (2019) [15] formally analyze 5G-AKA and show that it rests on unstated channel assumptions and suffers from session confusion. We do not study authentication; in TM2 we misuse the Control Plane on N4. By flooding PFCP Session Deletion Requests, we break existing User Plane sessions and trigger UAV failsafe, assuming that N4 is poorly segmented or protected. Salazar et al. (2021) [31] provide a 5G traffic fuzzer that can inject attacks on control and User Planes. We implement one concrete insider attack on N4 and tie it to the loss of the UAV-GCS session and the resulting failsafe behavior.
General drone vulnerabilities. Nassi et al. (2021) [26] systematize security and privacy for drones and for society; they do not focus on the cellular leg of the link. Branco et al. (2025) [12] review cyber attacks on commercial drones, including communication and C2. Work on the C2 protocol itself includes Du et al. (2024) [18], who exploit MAVLink weaknesses to hijack UAVs, and Veksler et al. (2024) [39], who use MAVLink for covert channels and information leakage; both operate at the application layer and assume direct or wireless access to the C2 channel. We instead assume the C2 channel is carried over 5G. We show how three 5G SA attacker positions (rogue UE in the same slice/DNN, insider on N4, compromised gNodeB) allow an attacker to affect that channel or the session that carries it, and we give testbed evidence and mitigations. In TM1 and TM3, MAVLink is the application protocol we attack; we recommend MAVLink signing so that injection and tampering by a rogue UE or a compromised gNodeB are rejected even when the 5G path is weak.
Table II summarizes how the cited works compare to this paper across seven criteria: focus on UE as attack vector or target, RAN (e.g., gNodeB), 5G core (e.g., N4/PFCP), User Plane (N3/GTP-U), Control Plane (N4/AKA or signaling), UAV or drone scenario, and testbed or empirical evaluation. A filled circle ($\bullet$) indicates that the criterion is a main focus or is explicitly addressed; an empty circle ($\circ$) indicates absence; a diamond ($\diamond$) indicates limited or partial presence. This work is the only one addressing all seven dimensions simultaneously in a controlled experimental testbed.